OpenStudy (anonymous):

How do I solve this system of equations algebraically? (Will give medal & fan) 10x + 4y = 24 -6x - 4y = 8

OpenStudy (carniel):

If I'm correct... your suppose to cancel y out 1st since it's easier to get x. so it will be 4x=-16

OpenStudy (carniel):

Divide it by 4 and you will get x which will be x=-4

OpenStudy (carniel):

Plug in x for any of those equations then subtract it from the sum

OpenStudy (carniel):

Like so, -6 x 4 = -24-4y=8

OpenStudy (carniel):

24+8 = 32

OpenStudy (anonymous):

Wait, the most we can do is just move y to the other side, right? we can't just get rid of it. so it would either be 10x = 24 - 4y for the first one or -6x = 8 + 4y for the second one.. right?

OpenStudy (carniel):

Divide by -4 which would be -8

OpenStudy (carniel):

You have to get rid of one to find the 1st one

OpenStudy (anonymous):

@phi

OpenStudy (carniel):

You can only do it if they cancel out. Which is why y was canceled out.

OpenStudy (anonymous):

I just don't understand...

OpenStudy (carniel):

y= -8 x=-4

OpenStudy (carniel):

I don't know how to explain it better than that :\ Sorry if I wasn't any help

OpenStudy (anonymous):

I graphed the system on a graphing calculator, so I know the solution is x = 8 and y = -14, I just don't know how to get it by doing the math

OpenStudy (carniel):

.... Well idk

OpenStudy (anonymous):

10(8) + 4(-14) = 24 80 + (-56) = 24 24 = 24

OpenStudy (anonymous):

-6(8) - 4(-14) = 8 -48 - (-56) = 8 8 = 8

OpenStudy (anonymous):

I just still need to be shown how to solve it algebraically... as if i never graphed it.. for anyone who wants to help...

OpenStudy (anonymous):

I don't know how to solve it algebraically, but I graphed it and the answer is (8, -14)

OpenStudy (anonymous):

Lol, I just saw that you knew that already...

OpenStudy (anonymous):

10x + 4y = 24 -6x - 4y = 8 You would first cancel out the 4y's to solve for x, then add the equations together. You would then have the equation 4x=32.

OpenStudy (anonymous):

Once you have that you would then divide both sides by 4.

OpenStudy (anonymous):

how do you cancel out the ys?

OpenStudy (anonymous):

Then you take that number and one of the equations and plug it in for x.

OpenStudy (anonymous):

Its simply adding 4y to -4y which equals 0

OpenStudy (anonymous):

So then it would be -6 = 8 + 4y though, right?

OpenStudy (anonymous):

you can't just add it to one side...

OpenStudy (anonymous):

No, you want to get rid of one to solve for the other.

OpenStudy (anonymous):

But... you can't just do something to one side and not the other.

OpenStudy (anonymous):

Yes you can, its called elimination.

OpenStudy (anonymous):

Let me show you....

OpenStudy (anonymous):

I've always been taught that you can't do something to one side and not the other, not in algebra.

OpenStudy (anonymous):

If you add 4y to one side, you have to add 4y to the other side of the equation as well.

OpenStudy (anonymous):

10x + 4y = 24 -6x - 4y = 8 4y+(-4y)=0 10x=24 + -6x=8 4x=32 X=8

OpenStudy (anonymous):

Your eliminating so you can solve for x then you are going to use that answer to solve for y.

OpenStudy (anonymous):

I just don't even.... I was never taught anything like that.

OpenStudy (anonymous):

It will be much easier once you learn that. Then things become less confusing.

OpenStudy (anonymous):

I just don't know if I'll get credit for showing work I was never taught...

OpenStudy (anonymous):

But I do see how it works... sorta

OpenStudy (anonymous):

You cant solve it by moving them to different sides.

OpenStudy (anonymous):

Its literally impossible to do that.

OpenStudy (anonymous):

Oh okay, I see now, I went through my notes and found something similar

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Your welcome! :)

OpenStudy (phi):

I am sure they taught you two ways to solve these equations one way is substitution. The other is elimination. elimination (in my opinion) is easier, especially if you have more than 2 equations and 2 unknowns.

OpenStudy (anonymous):

I already figured this problem out, I forgot about the other method but then found something similar in my notes after a little bit of digging. I do need help on a different problem now though.

OpenStudy (phi):

10x + 4y = 24 -6x - 4y = 8 in this problem, right away I notice a +4y in the first equation and a -4y in the 2nd equation... that is a clue to use elimination. Here is the idea: the left side of eq 1 is equal to 24 the left side of equation 2 is equal to 8 if we add the left sides of the equations, the sum will be 8+24= 32 so let's do that: 10x + 4y + -6x + -4y= 32 (we are doing this because we know 4y + -4y is 0, and y is gone!) 10x - 6x simplifies to 4x and we get 4x = 32 x= 8 now use either of the equations to find y